### v1

@@ Abstract
Introduction
Over the last decade, researchers in quantum information processing have broadly divided their efforts into two distinct but complementary themes. In one, the focus has been on realizing the building blocks for large-scale, fault-tolerant quantum processors [28, 24, 7], which would enable running algorithms such as Shor’s or Grover’s at meaningful scale. In another, there has been a push to realize quantum systems comprising tens to hundreds of qubits or qumodes, but without error correction, and to explore what can be done with such noisy, pre-fault-tolerance systems—often denoted as noisy, intermediate-scale, quantum (NISQ) devices [38]. Quantum computational supremacy with such NISQ devices has been demonstrated [2, 49], but there has been much less progress on achieving quantum advantage in practically relevant applications than had been hoped for as NISQ machines began to be created [35]. There have been many NISQ studies on quantum machine learning [5], and in this area too, quantum advantage for problems of broad practical interest has remained elusive [40, 8]. A major open question is whether one can achieve any practically relevant advantage for machine learning with NISQ systems.
One of the main approaches to performing quantum machine learning with NISQ machines is to use parameterized quantum circuits as quantum neural networks [42, 25, 9], which are a subclass of variational quantum algorithms, in which parameters of a quantum circuit are adjusted, usually by a classical co-processor, so that the quantum circuit incrementally approaches carrying out a desired computation. This approach, however, typically suffers from barren plateaus [32, 44, 31, 1], which mean that, in practice, it is difficult or impossible to perform the optimization required to set circuit parameters [19]. Inspired by the framework of reservoir computing [29, 39, 26, 17] in classical machine learning, quantum reservoir computing (QRC) [15, 14, 18] has emerged as an approach to quantum machine learning that entirely avoids barren plateaus by performing all the learning in the classical domain. They key idea of a QRC is that a quantum system (called a quantum reservoir) can generate nonlinear, high-dimensional features of inputs to it, and that these features can be used to perform machine-learning tasks purely by training a classical linear transformation. QRC can be implemented both in the circuit model of quantum computation [14] and with analog quantum dynamical systems [30, 6]. However, experimental demonstrations to date have been performed with digital quantum circuits [37, 10, 27, 34, 48, 41, 21] that have limited the complexity of tasks that can be performed, in part due to an input bottleneck imposed by the use of discrete gates to input temporal data using a series of separate, imperfect gates.
The aim of our work is to demonstrate a proof-of-principle for a new application of and approach to quantum machine learning with NISQ devices that overcomes or sidesteps the challenges in training and inputs noted above. We use the driven, continuous-time analog quantum nonlinear dynamics of a superconducting microwave circuit as a quantum reservoir to generate features for classifying weak, analog microwave signals (Fig. 1a). We use repeated measurements of the reservoir both to extract features that contain information about temporal correlations in the input data, as well as to induce non-unitary dynamics. Our use of a continuous-variable system in our quantum reservoir grants us access to a substantially larger Hilbert space than would be the case with a qubit-only system with equally many hardware components. In relying on continuous-time dynamics, our approach is similar to other proposals for analog NISQ processors and simulators [36, 16, 11], which aim to avoid the overhead that imposing a discrete-time (circuit-model, gate-based) abstraction causes. Analog operation grants us an even more important ability however, which fundamentally distinguishes our work from prior experimental demonstrations of quantum machine learning on circuit-model quantum processors: it allows our device to directly, natively receive weak analog microwave signals, and to immediately leverage analog quantum information processing to extract relevant features of the signals for classification.
This small shift in context has important implications, offering a new path to practical quantum advantage with NISQ hardware. Rather than focusing on using NISQ hardware to perform computation on pre-recorded, digital data, we instead use quantum hardware to perform computation on real-time analog signals that interface directly with our microwave superconducting device. Our experiments do not address the question of whether a QRC can achieve a quantum computational advantage, since our experimental device is small enough to be easily classically simulable. However, our demonstrations suggest a route to achieving a quantum advantage of a different kind: an advantage in the quantum detection and processing of weak microwave signals, allowing quantum hardware to extract complex information of interest from dim, analog signals in ways that would be noisier with a conventional classical approach. This type of quantum advantage, arising from a combination of quantum sensing with extraction of complex features about the sensed signal, is discussed in general terms as a route to quantum advantage with quantum machine learning in Ref. [8]. Our work shows that when classical signals comprising just a few photons have entered an analog quantum reservoir, they can be classified using our QRC approach. If one combines this analog quantum processing with a sensitive quantum detector of microwave radiation, as has already been previously demonstrated using superconducting circuits [46, 45, 4, 3], then one can construct a system that achieves a quantum advantage in the task of combined sensing and signal processing.
Figure 1: Analog signal classification with a continuous-variable quantum reservoir computer (QRC) using measurement trajectories. (a.) We perform machine learning using a quantum system consisting of an oscillator coupled to a qubit. Analog signals are fed into our analog quantum dynamical system, continually displacing the oscillator mode while the qubit is protectively measured. The measurement trajectories provide complex features a digital linear layer can use to perform classification on a variety of tasks. (b.) The signals interface directly with the qubit-oscillator system, composed of a 3D aluminum cavity hosting a transmon qubit. (c.) Simulated Wigner distributions of the oscillator state through various stages of the reservoir dynamics. The dynamics include entanglement-generating unitary evolution, and projective measurements of both the qubit and cavity. The back-action produced by the measurements add to the non-classical features generated by the entangling unitaries. The balance of measurements and unitaries, which do not commute with each other in our implementation, lead to complex correlations in the measurement trajectories. (b.) The digital linear layer performs classification based on a feature vector, which we construct using the expectation values of the central moments $\mu_{p}$, which capture the essential correlations in the reservoir dynamics.
@@ Introduction
Figure 2: Reservoir protocol overview with an example time-independent classification task (a) The unitary dynamics in our reservoir is generated by control pulses that serve to entangle the qubit with the oscillator before the analog input is received by the oscillator. For tasks where the analog data is time-independent, the dynamics is fully gate-based, and the oscillator is dis-entangled with the qubit before the qubit and oscillator measurements. For details of the motivation behind the particular unitaries implemented for our reservoir, see Appendix C. (b) (Inset) An illustrative machine learning example is the classification of time-independent signals from two arms of a Spiral distribution defined in the signal $I-Q$ plane. Unlike classical machine learning, in quantum machine learning, the performance is unavoidably impacted by sampling noise. Here, we plot the classification accuracy of the spiral task with increasing number of shots. Also plotted is the performance of a linear layer acting directly on the two-dimensional $I,Q$ data, indicating that non-linearity is required to perform this task with sufficient accuracy. (c) Classification accuracy at $10^{4}$ shots as a function of qubit coherence time that we tune via resonator-induced dephasing during the classification (see Appendix D). While we see a large drop in classification performance when the qubit coherence time is heavily suppressed and the system is completely disentangled, the performance only begins to suffer once the qubit $T_{2}$ approaches the duration between measurements.
Experimental setup and protocols
Our quantum reservoir, composed of a long-lived cavity mode coupled to a transmon qubit (Fig. 1b), can be modeled with the rotating-frame Hamiltonian,
$H=-\chi a^{\dagger}a\frac{\sigma_{z}}{2}+\epsilon^{*}(t)a+\epsilon(t)a^{\dagger}+\Omega_{x}(t)\sigma_{x}+\Omega_{y}(t)\sigma_{y},$ (1)
where $\sigma_{z}$ is the Pauli operator on the qubit subspace of the transmon, $a$ is the photon annihilation operator of the cavity mode, and $\chi$ is the nonlinear interaction strength (see Appendix C for details). The right-most term of Eq. 1 describes the unitary control of the qubit, and the second term describes both the encoding of the input data $\epsilon_{\mathrm{in}}(t)$, and unitary control of the oscillator mode, i.e., $\epsilon(t)=\epsilon_{\mathrm{in}}(t)+\epsilon_{\mathrm{control}}(t)$. Equation 1 describes the unitary dynamics, which is complemented by non-unitary dynamics generated by the back-action from qubit measurements interspersed throughout the evolution.
The oscillator and qubit control drives used in our work realize a reservoir that consists of a series of entangling unitaries interleaved with qubit and oscillator measurements (Fig. 1c). The analog input results in a time varying displacement of the cavity, which streams in concurrently with control drives implementing an entangling unitary. Following the unitary, we perform a qubit measurement, and then the parity of the oscillator state is measured [43, 20] (see Appendix D). The parity measurement projects the oscillator state into either even or odd super-positions of Fock states, giving us sensitive information about photon number changes of oscillator while inducing non-classical features to the state via measurement back-action. In effect, our construction implements a sequence of non-commuting measurements (see Appendix C), generating correlated measurement distributions that can then be used as complex output features.
@@
The measurement outcomes are used to construct output feature vectors to be fed into the linear layer (Fig. 1a), but this can be done in a few different ways. In principle, with repeated applications of the unitary, we generate a sample of bitstrings with $2^{M}$ possible outcomes, where $M$ is the number of measurements. The outcomes can be counted to directly form a sample probability distribution over measurement trajectories, which can then be used as a high-dimensional output feature vector after obtaining a sufficient number of samples $N$. While this approach has the benefit of capturing all information in the measurement distribution [21], it can generally suffer from poor scaling in sampling noise, requiring $N\sim 2^{M}$ shots in the worst case [47]. On the other hand, one could average over the measurements directly [48, 41]; however, this has the unwanted effect of averaging over and removing quantum correlations. Here, we construct an output feature vector from estimates of successive central moments $\mu_{1},\mu_{2},\mu_{3},\ldots$ of the underlying distribution over measurement trajectories (Fig. 1d). Additionally, given finite memory in our reservoir, we choose to only use correlations between measurements at most 3 measurements apart. This approach, inspired by Ref. [26], has the benefit of leveraging the hierarchy of noise in the central moments, while capturing the essential correlations in the dynamics to achieve high accuracy even in the few-sample regime. For a detailed analysis of the construction of our reservoir output features with comparisons, see Appendix E.
@@ Results
To illustrate the scheme proposed in this work, we begin with an example classification using our quantum reservoir by performing binary classification task of time-independent signals. Figure 2a describes the control drives in more detail. For time-independent input data, the two-dimensional input data is encoded as the $I$ and $Q$ quadratures of an analog signal resonant with the oscillator frequency, which displaces the cavity. Here, time-independence describes the signal refers to the fact that it is on resonance the oscillator mode, and thus has no time-dependence in the oscillator’s rotating frame (such that $\epsilon_{\mathrm{in}}(t)=\epsilon_{\mathrm{in}}$ in Eq. 1). For time-independent tasks, the signal bandwidth is set by its duration, and therefore the resultant displacement is essentially conditioned on the qubit in the ground state due to the cross-Kerr interaction.
@@
The unitary encoding the input displacement is complimented by control drives that entangle the qubit and cavity via a series of conditional displacements [13] and qubit rotations. For time-indepedent tasks, this set of unitaries effectively impart a geometric area enclosed by the cavity trajectory onto the qubit, such that the information of the phase of the unknown input signal can be extracted via a qubit measurement (see Appendix C for details of this unitary). In Appendix H, we show the ability of the set of unitaries implemented here (Fig. 2a) to be able to approximate any scalar function of the input signal when the signal is time-independent. For all results presented, we implement our reservoir unitary with these control drives across all tasks, with 4 applications of the unitary interleaved with qubit and oscillator-parity measurements.
The binary classification task we perform here is: Two distributions of time-independent signals, completely characterized by the signal’s in-phase ($I$) and quadrature ($Q$) components, are distributed along two separate “arms of a spiral” in the $I-Q$ plane (Fig. 2b). The classification task is: given a displacement described by the points $I$ and $Q$ sampled from either signal distribution, figure out which distribution the signal came from. This simple task has the feature that, if one feeds in the inputs directly into a linear layer, this would classify with an accuracy of no more than $67\%$ – just above random guessing of $50\%$ (Fig. 2b). As a point of comparison with non-linear digital networks, we found that a 64-dimensional, two-layer digital reservoir was needed to achieve the same performance as our quantum reservoir for this task (see Appendix J for details of this comparison).
@@
To probe the role of quantum in our reservoir, we performed the same classification task, but with reduced coherence time in the qubit during the reservoir execution. This is achieved by populating the lossy readout resonator with photons that send the qubit to the center of the Bloch-sphere when the readout resonator is traced out (see Appendix D). With $T_{2}\rightarrow 0$, we effectively removed all entanglement with the cavity, and observed two things: a dramatic reduction in classification performance, and importantly, $T_{2}$ only began affecting the performance once it was on the order of the reservoir duration, after which the qubit was projected. This latter point highlights an important benefit of repeated measurements in our reservoir construction, i.e. while entanglement is important for generating complex distributions in our setup, we are able to classify and capture information much longer than the qubit decoherence time, requiring only that the oscillator state is coherent at long times.
@@
We showcase our reservoir in a real-world setting by discriminating time-dependent radio-frequency (RF) signals from 10 different digital modulation schemes. Digital modulation schemes encode binary information in discrete ‘symbols’ encoding in sequential time-bins. For example, Binary Phase-Shift Keying (BPSK) encodes binary data in discrete phase jumps of a signal, such that a symbol 00 ($1$) maps to a phase flip of 00 ($\pi$). While BPSK only contains one bit of information per symbol, other encoding schemes such as 32 Quadrature Amplitude Modulation (32QAM) can encode 5 bits per symbol. These and other encodings can be represented in a constellation diagram (Fig. 3a), which denotes the potential $(I,Q)$ values a signal can take for each symbol. A given string of digital data can then be encoding in a time-domain signal by sequentially choosing points in the constellation diagram with a given symbol rate, denoting the rate at which the symbol will change. For typical WiFi signals this is around 250 kHz per subchannel [23].
@@
Figure 3: Classification of radio-frequency (RF) communication modulation protocols. (a) Description of the dataset for digital modulation schemes used in this experiment. In conventional digital modulation schemes, one encodes data in the amplitude and phase of the signal. The modulation schemes can be represented by a “constellation diagram” in $(I,Q)$ space (left), where point represents one of the possible choice of $(I,Q)$ values to enclode a symbol, with example time traces (right). (b) These signals are broken up and fed into our reservoir. (c) The performance of the reservoir as a function of the shots taken in real time (see text). The top row contains the corresponding duration of the radio frequency signal required. As the number of shots increases, the fluctuations in the measurement distribution reduces, resulting in a higher classification accuracy. For context, a classical linear classifier applied directly on the input data achieves only $20\%$ accuracy, independent of the duration of the signal. The error bars of the accuracy curve have been omitted here due to the fact they are too small. (d) Confusion matrix for the QRC at 32, 512, and $10^{4}$ shots, showing that the reservoir confuses only a few classes at the highest shots.
Figure 4: Classification of filtered noise (a) We classify various fast and slow noisy signals by applying a moving average on stochastic white-noise signals. Three different filters are used for the moving averages: a Gaussian filter, a Lorentzian, and an inverse power law. For each filter, we generate stochastic analog signals based on both a 50 ns filter width, and a 600 ns filter width, with the latter being on the order of the measurement rate. Example time traces are given for the real part of an example stochastic white noise signal passed through each of the filters. We divide up these stochastic signals and send to the QRC to then distinguish the noise source. (b) Visualization of the high-dimensional output feature space using Singular-Value-Decomposition. Each point corresponds to a different signal over 2000 shots taken in real time (see text). (c) Classification accuracy as a function of the number of shots using third order moments as the output feature. Our reservoir reaches 93% accuracy in about 2000 shots, corresponding to about 10 ms of the signal received. (d) Confusion matrix of the task taken at 2000 shots. (e) Participation of the mean and the off-diagonal elements of the second- and third-order moments in the classification accuracy within the subtasks of classifying different noise sources with fixed filter width. We see that for signals with long coherence times, higher order measurement correlations are important, while for fast signals, the mean dominates in the performance contribution.
@@
For this task, we generated RF signals by encoding random digital strings into the 10 different modulation schemes with a fixed symbol rate of around 2 symbols per $\mu$s. The duration of these signals can last much longer than the reset rate of our system. Importantly, we did not repeat the same signal to artificially reduce the sampling noise associated with each input data, as this would not typically be applicable in a real-world setting. Instead, the measurement statistics were generated by sampling the signal in real time. Consequently, what we refer to as ‘shots’ in a real-time task do not correspond to identical repetitions of the experiment, but instead, is the number of resets we performed while acquiring the signal, which changed from shot to shot. In effect, each different encoding scheme produces a unique “fingerprint” of distribution over measurement outcomes, and it is the goal of the linear layer is to separate these distributions with as high accuracy as possible.
@@
Figure 3c shows the accuracy in classifying digitally modulated RF signals with increasing number of shots, compared with the performance of a linear classifier. We note that in less than a millisecond, or with less than 2000 symbols, the reservoir was able to classify which of the 10 classes a given signal belongs to with $>$ 90% accuracy when using 8 qubit-cavity measurements. A linear classifier can only achieve 20% classification accuracy for this task, even with infinite symbols. The confusion matrix between the difference classes at 32, 512, and $10^{4}$ shots is displayed in Fig 3d, which is nearly diagonal.

### v2

@@ Abstract
Quantum reservoir computing (QRC) has been proposed as a paradigm for performing machine learning with quantum processors where the training is efficient in the number of required runs of the quantum processor and takes place in the classical domain, avoiding the issue of barren plateaus in parameterized-circuit quantum neural networks. It is natural to consider using a quantum processor based on superconducting circuits to classify microwave signals that are analog—continuous in time. However, while theoretical proposals of analog QRC exist, to date QRC has been implemented using circuit-model quantum systems—imposing a discretization of the incoming signal in time, with each time point input by executing a gate operation. In this paper we show how a quantum superconducting circuit comprising an oscillator coupled to a qubit can be used as an analog quantum reservoir for a variety of classification tasks, achieving high accuracy on all of them. Our quantum system was operated without artificially discretizing the input data, directly taking in microwave signals. Our work does not attempt to address the question of whether QRCs could provide a quantum computational advantage in classifying pre-recorded classical signals. However, beyond illustrating that sophisticated tasks can be performed with a modest-size quantum system and inexpensive training, our work opens up the possibility of achieving a different kind of advantage than a purely computational advantage: superconducting circuits can act as extremely sensitive detectors of microwave photons; our work demonstrates processing of ultra-low-power microwave signals in our superconducting circuit, and by combining sensitive detection with QRC processing within the same system, one could achieve a quantum sensing-computational advantage, i.e., an advantage in the overall analysis of microwave signals comprising just a few photons.
Introduction
Over the last decade, researchers in quantum information processing have broadly divided their efforts into two distinct but complementary directions. In one, the focus has been on realizing the building blocks for large-scale, fault-tolerant quantum processors [27, 23, 6], which would enable running algorithms such as Shor’s or Grover’s at meaningful scale. In the other, there has been a push to realize quantum systems comprising tens to hundreds of qubits or qumodes, but without error correction, and to explore what can be done with such noisy, pre-fault-tolerance systems—often denoted as noisy, intermediate-scale, quantum (NISQ) devices [38]. Quantum computational supremacy with such NISQ devices has been demonstrated [2, 50], but there has been much less progress on achieving quantum advantage in practically relevant applications [34]. There have been many NISQ studies on quantum machine learning [5], and in this area too, quantum advantage for problems of broad practical interest has remained elusive [41, 7]. A key challenge in quantum neural networks realized with parameterized quantum circuits has been training the parameters when the optimization landscape suffers from barren plateaus [31, 45, 29, 1]. A major open question is whether one can achieve any practically relevant advantage for machine learning with NISQ systems.
Figure 1: Analog signal classification with a continuous-variable quantum reservoir computer (QRC) using measurement trajectories. (a.) We perform machine learning using a quantum system consisting of an oscillator coupled to a qubit. Analog signals are fed into our analog quantum dynamical system, continually displacing the oscillator mode while the qubit is projectively measured. The measurement trajectories provide complex features a digital linear layer can use to perform classification on a variety of tasks. (b.) The signals interface directly with the qubit-oscillator system, composed of a 3D aluminum cavity (blue) hosting a transmon qubit (red) and a readout resonator (omitted for clarity; see Appendix A). (c.) Wigner tomography performed on the oscillator state through various stages of the reservoir dynamics. The dynamics include entanglement-generating unitary evolution, and projective measurements of both the qubit and oscillator. The back-action produced by the measurements add to the non-classical features generated by the entangling unitaries. The balance of measurements and unitaries, which do not commute with each other in our implementation, lead to complex correlations in the measurement trajectories. (d.) The digital linear layer performs classification based on a feature vector, which we construct using the expectation values of the central moments $\mu_{p}$ ($p=1,2,3,\ldots$), which capture the essential correlations in the reservoir dynamics.
Inspired by the framework of reservoir computing [28, 40, 25, 15] in classical machine learning, quantum reservoir computing (QRC) [13, 12, 16, 17, 24, 42] has emerged as an approach to quantum machine learning that entirely avoids barren plateaus by performing all learning in a final, linear layer. They key idea of a QRC is that a quantum system (called a quantum reservoir) can generate nonlinear, high-dimensional features of inputs to it, and that these features can be used to perform machine-learning tasks purely by training a classical linear transformation. However, experimental demonstrations to date have been performed with digital quantum circuits [37, 8, 26, 33, 49, 43, 19] that have limited the complexity of tasks that can be performed, in part due to an input bottleneck imposed by the use of discrete gates to input temporal data using a series of separate, imperfect gates.
The aim of our work is to demonstrate a proof-of-principle for a new application of and approach to quantum machine learning with NISQ devices that overcomes or sidesteps the challenges in training and inputs noted above. We use the driven, continuous-time analog quantum nonlinear dynamics of a superconducting microwave circuit as a quantum reservoir to generate features for classifying weak, analog microwave signals (Fig. 1a). We use repeated measurements of the reservoir both to extract features that contain information about temporal correlations in the input data, as well as to induce non-unitary dynamics. Our use of a continuous-variable system in our quantum reservoir grants us access to a substantially larger Hilbert space than would be the case with a qubit-only system with equally many hardware components. Our approach is similar to proposals for analog NISQ processors and simulators [36, 14, 9], which aim to avoid the overhead caused by imposing a discrete-time abstraction. Analog operation grants us an even more important ability however, which fundamentally distinguishes our work from prior experimental demonstrations of quantum machine learning on circuit-model quantum processors: it allows our device to directly, natively receive weak analog microwave signals, and to immediately leverage analog quantum information processing to extract relevant features of the signals for classification.
Rather than focusing on using NISQ hardware to perform computation on pre-recorded, digital data, we instead use quantum hardware to perform computation on real-time analog signals that interface directly with our microwave superconducting device. Our experiments do not address the question of whether a QRC can achieve a quantum computational advantage, since our experimental device is small enough to be easily classically simulable. However, our demonstrations suggest a route to achieving a quantum advantage of a different kind: an advantage in the quantum detection and processing of weak microwave signals, allowing quantum hardware to extract complex information of interest from dim, analog signals in ways that would be noisier with a conventional classical approach. This type of quantum advantage, arising from a combination of quantum sensing with extraction of complex features about the sensed signal, is discussed in general terms as a route to quantum advantage with quantum machine learning in Ref. [7]. Our work shows that when classical signals comprising just a few photons have entered an analog quantum reservoir, they can be classified using our QRC approach. The signals we classify are synthesized at room temperature and pass through 60 dB of attenuation before reaching our device. However, if instead one combines this analog quantum processing with a sensitive quantum detector of microwave radiation, as has already been previously demonstrated using superconducting circuits [47, 46, 4, 3, 11], then one can construct a system that achieves a quantum advantage in the task of combined sensing and signal processing of high temperature signals.
Experimental setup and protocols
Our quantum reservoir, composed of a cavity resonator coupled to a transmon (Fig. 1b), can be modeled with the following qubit-oscillator Hamiltonian in the rotating-frame,
$H/\hbar=-\chi|e\rangle\langle e|a^{\dagger}a+\epsilon(t)a^{\dagger}+\Omega(t)|e\rangle\langle g|+\mathrm{H.c.},$ (1)
where $|g\rangle$ and $|e\rangle$ define the qubit subspace of the transmon, $a$ is the photon annihilation operator of the oscillator mode, and $\chi$ is the nonlinear interaction strength (see Appendix B for details). The third term of Eq. 1 describes the unitary control of the qubit with a time-dependent drive $\Omega(t)$, and the second term describes both the encoding of the input data $\epsilon_{\mathrm{in}}(t)$, and unitary control of the oscillator mode, i.e., $\epsilon(t)=\epsilon_{\mathrm{in}}(t)+\epsilon_{\mathrm{control}}(t)$. Equation 1 describes the unitary dynamics, which is complemented by non-unitary dynamics generated by the back-action from qubit measurements interspersed throughout the evolution.
@@ Introduction
Figure 2: Reservoir protocol overview with an example time-independent classification task (a) The unitary dynamics in our reservoir are generated by control pulses that serve to entangle the qubit with the oscillator before the analog input is received by the oscillator. For tasks where the analog data is time-independent, the dynamics are fully gate-based, and the oscillator is dis-entangled with the qubit before the qubit and oscillator measurements. For details of the motivation behind the particular unitaries implemented for our reservoir, see Appendix B. (b) (Inset) An illustrative machine learning example is the classification of time-independent signals from two arms of a Spiral distribution defined in the signal $I-Q$ plane. For quantum machine learning, unlike classical, the performance is unavoidably impacted by sampling noise. Here, we plot the classification accuracy of the spiral task against an increasing number of shots. Also plotted is the performance of a linear layer acting directly on the two-dimensional $I,Q$ data, indicating that non-linearity is required to perform this task with sufficient accuracy. (c) Classification accuracy at $10^{4}$ shots as a function of qubit coherence time that we tune via resonator-induced dephasing during the classification (see Appendix C). While we see a large drop in classification performance when the qubit coherence time is heavily suppressed and the system is completely disentangled, the performance only begins to suffer once the qubit $T_{2}$ approaches the duration between measurements.
@@
The measurement outcomes are used to construct output feature vectors to be fed into the linear layer, but this can be done in a few different ways. When performing repeated measurements on our system, we generate a sample bitstring of length $M$ describing the quantum trajectory over $M$ measurements. After $M$ measurements are performed, we reset the system and repeat the procedure, each time generating a bitstring $\vec{x}_{n}=[x_{n0},x_{n1},\ldots,x_{nM-1}]$, where $n$ refers to the $n$th sample (Fig. 1c). The outcomes can be counted to directly form a sample probability distribution $p(\vec{x}|\epsilon_{\mathrm{in}}(t))$ over measurement trajectories, which can then be used as a high-dimensional output feature vector after obtaining a sufficient number of samples $N$. While this approach has the benefit of capturing all information in the measurement distribution [19], it can generally suffer from poor scaling in sampling noise, requiring $N\sim 2^{M}$ shots in the worst case [48].
Here, we construct an output feature vector from estimates of successive central moments $\mu_{1},\mu_{2},\mu_{3},\ldots$ of the underlying distribution $p(\vec{x}|\epsilon_{\mathrm{in}}(t))$ (Fig. 1d). For example, the first-order central moment $\mu_{1}$ is a $M$-dimensional vector representing the average over all measured bitstrings, i.e. $\mu_{1}=[\langle x_{n0}\rangle,\langle x_{n1}\rangle,\ldots]$, the second-order central moment $\mu_{2}$ is the covariance matrix with elements $(\mu_{2})_{ij}=\langle x_{ni}x_{nj}\rangle-\langle x_{ni}\rangle\langle x_{nj}\rangle$, and so on. Here, the expectation value is taken over the sample index $n$. This approach, inspired by Ref. [25], has the benefit of leveraging the hierarchy of noise in the central moments, while capturing the essential correlations in the dynamics to achieve high accuracy even in the few-sample regime. Furthermore, the output feature vector dimension only scales polynomially with the number of measurements, where the highest polynomial power is given by the order of the highest central moment, which we restrict to 3 for all tasks in this work. Finally, given finite memory in our reservoir, we further restrict the output vector by choosing to only calculate correlations between measurements at most 3 measurements apart. These truncated moments are then flattened and concatenated to construct our output feature vectors. In all, for the $M$ = 8 measurements we use in this work, the resultant output feature vector size with this prescription is 94. For a detailed discussion of the construction of our reservoir output features with comparisons of different encodings, see Appendix D.
@@ Results
To illustrate the scheme proposed in this work, we begin with an example classification using our quantum reservoir by performing binary classification task of time-independent signals. Figure 2a describes the control drives in more detail. For time-independent input data, the two-dimensional input data is encoded as the $I$ and $Q$ quadratures of an analog signal resonant with the cavity resonance frequency. In the rotating frame of the system (Eq. 1), this is effectively a time-independent signal, i.e. $\epsilon_{\mathrm{in}}(t)=\epsilon_{\mathrm{in}}=I+Q$, which results in a displacement of the oscillator state. For such time-independent tasks, the signal bandwidth is set by its duration which, in general, can make the resultant displacement conditioned on the qubit state due to the cross-Kerr interaction (see first term of Eq. 1).
@@
The unitary encoding the input displacement is complimented by control drives that entangle the qubit and oscillator via conditional displacements [10] and qubit rotations (Fig. 2a). The entangling conditional displacements are applied before and after the unknown input is fed into the system, and the qubit is rotated by $\pi$ or $\pi/2$ pulses before, during, and after the input. Due to the qubit-state-dependent shift of the oscillator frequency by $-\chi$, these qubit rotations serve to make the oscillator sensitive to the input signal independent of the state of the qubit at the start of each round of input. Additionally, when combined with conditional displacements on the oscillator, the control and input scheme impart a geometric area enclosed by the oscillator trajectory onto the qubit, such that the phase of an unknown time-independent input signal can be extracted via a qubit measurement (see Appendix B for details of this unitary). In Appendix G, we show the ability of the set of unitaries implemented here to be able to approximate any scalar function of the input signal when the signal is time-independent. For all results presented, we implement our reservoir unitary with these control drives across all tasks, with 4 applications of the unitary interleaved with qubit and oscillator-parity measurements.
[…]
@@
To probe the role of quantum coherence in our reservoir, we performed the same classification task, but with reduced coherence time in the qubit during the reservoir execution (Fig. 2c). This was achieved by populating the lossy readout resonator with photons that send the qubit to the center of the Bloch-sphere when the readout resonator is traced out (see Appendix C). With $T_{2}\rightarrow 0$, we effectively removed all entanglement with the oscillator, and observed two things: a dramatic reduction in classification performance, and importantly, $T_{2}$ only began affecting the performance once it was on the order of the reservoir duration, after which the qubit is projected to a pure state.
@@
Figure 3: Classification of radio-frequency (RF) communication modulation protocols. (a) Description of the dataset for digital modulation schemes used in this experiment. In conventional digital modulation schemes, one encodes data in the amplitude and phase of the signal. The modulation schemes can be represented by a “constellation diagram” in $(I,Q)$ space (left), where point represents one of the possible choice of $(I,Q)$ values to encode a symbol, with example time traces (right). (b) These signals are broken up and fed into our reservoir. (c) The performance of the reservoir as a function of the shots taken in real time (see text). The top row contains the corresponding duration of the radio frequency signal required. As the number of shots increases, the fluctuations in the measurement distribution reduces, resulting in a higher classification accuracy. For context, a classical linear classifier applied directly on the input data achieves only $20\%$ accuracy, independent of the duration of the signal. The error bars of the accuracy curve have been omitted here due to the fact they are too small. (d) Confusion matrix for the QRC at 32, 512, and $10^{4}$ shots, showing that the reservoir confuses only a few classes at the highest shots.
Next, to highlight the ability to perform classification of higher dimensional data, we classified time-dependent radio-frequency (RF) signals. The microwave signals in this dataset encode digital information using one of 10 different digital modulation schemes, a standard benchmark task in RF machine learning [35, 21]. Digital modulation schemes encode binary information in discrete ‘symbols’ encoding in sequential time-bins. For example, Binary Phase-Shift Keying (BPSK) encodes binary data in discrete phase jumps of a signal, such that a symbol 00 ($1$) maps to a phase flip of 00 ($\pi$). Other modulation schemes can encode more bits per symbol. BPSK and other encodings can be represented in a constellation diagram (Fig. 3a), which denotes the potential $(I,Q)$ values a signal can take for each symbol. A given string of digital data can then be encoded in a time-domain signal by sequentially choosing points in the constellation diagram with a given symbol rate. For typical WiFi signals this is around 250 kHz per subchannel [22].
@@
For this task, we generated RF signals by encoding random digital strings into the 10 different modulation schemes with a fixed symbol rate of around 2 symbols per $\mu$s, or with a sampling rate of 2 MSps. The duration of these signals typically lasts much longer than the reset period of our system. Importantly, we did not repeat the same signal to artificially reduce the sampling noise associated with each input data, as this would not typically be applicable in a real-world setting. Instead, the measurement statistics were generated by sampling the signal in real time. Consequently, what we refer to as ‘shots’ in a real-time task does not correspond to identical repetitions of the experiment, but instead, is the number of resets we performed while acquiring the signal, which changed from shot to shot. In effect, each different encoding scheme produces a unique “fingerprint” distribution over measurement outcomes, and the goal of the linear layer is to separate these distributions with as high accuracy as possible.
@@
Figure 3c shows the accuracy in classifying digitally modulated RF signals with increasing number of shots, compared with the performance of a linear classifier. We note that in less than a millisecond, or with less than 2000 symbols, the reservoir was able to classify which of the 10 classes a given signal belongs to with $>$ 90% accuracy when using 8 qubit-oscillator measurements. A linear classifier can only achieve 20% classification accuracy for this task, even with infinite symbols. The confusion matrix between the different classes at 32, 512, and $10^{4}$ shots is displayed in Fig 3d, the latter two of which are nearly diagonal.